OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2? .will give medal and fan

OpenStudy (anonymous):

@aaronq @amistre64 @ganeshie8 @mathmale @Preetha

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Yalda6 @uri @tHe_FiZiCx99

OpenStudy (anonymous):

@Palmo4ka @pmuldowney @phi @Palmo4ka @amistre64 @mathmale @

OpenStudy (anonymous):

@johnweldon1993 @nincompoop @ajprincess @aaronq

OpenStudy (anonymous):

do you knowwhatexplicit form is

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wow somethings wrong with my space bar today 'know what explicit form is'

OpenStudy (anonymous):

explicit form is \[a _{n}=a1+(n-1)d\] a1= starting point n=what you're looking for (say i need the 25th number id put 25) d= the difference between numbers

OpenStudy (anonymous):

k let me substitute

OpenStudy (anonymous):

an=5+(n-1)-3

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

k dont help man

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you're expressing that you're subtracting everything in parenthesis by 3

OpenStudy (anonymous):

no its supposed to multiply everything in parenthesis by negative 3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

and how do we express that?

OpenStudy (anonymous):

parenthesis is a mathematicians best friend

OpenStudy (anonymous):

i don't know how you'd express it man

OpenStudy (anonymous):

can you do me a favor

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[x=(a)(-b)\] \[x=a-b\]

OpenStudy (anonymous):

can you tell me the difference between those equations

OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2? .will give medal and fan

OpenStudy (anonymous):

sorry i coudnt see the question so i copied it down again

OpenStudy (anonymous):

one of them is multiplication while the other is substraction

OpenStudy (anonymous):

yes and how did we change it from subtraction to multiplication

OpenStudy (anonymous):

parenthesis?

OpenStudy (anonymous):

yes so \[a _{n}=5+(n-1)-3\] this is saying we are subtracting by 3 how do we fix that into multiply by -3?

OpenStudy (anonymous):

(-3)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k so: an=5+(n-1)(-3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2? .will give medal and fan

OpenStudy (anonymous):

how do i find the domain

OpenStudy (anonymous):

and the explicit equation?

OpenStudy (anonymous):

we already did the explicit equation https://answers.yahoo.com/question/index?qid=20140224190617AAYdUcE read what they said on here and tell me how they found the domain

OpenStudy (anonymous):

k thnks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

@KlOwNlOvE